How does the compound C9H12O look based on this NMR graph?1: C3H120 S, 3H t, 3H Hz ppm 650.38 7.262 30 643.50 7.186 39 642.44 7.174 54 641.19 7.160 51 636.81 7.111 31 634.31 7.083 67
Answer #1

Molecular formula = Cg Hi 2o Degree of unsaturation = 26+2+ N-X-14 - 2x9+2-12 2 = 4 2 Multiplicity = 2n1+l, where I= !, & n=13C NMR: G → give singlelat 156 so bom G = give singlet at 29.01 tom I C=give singlet at 55.04ppm C4 + give singlet at 11.00